OpenStudy (anonymous):

Add any of the 5 numbers from 2, 6, 10, 14, 18, 22, 26, 30, 34, 38 and get the sum as 100.

OpenStudy (aum):

I am thinking it may not be possible.

OpenStudy (aum):

The numbers are in arithmetic sequence and there are 10 terms. The kth term is: \(2 + (k-1)d = 2 + (k - 1)4 = 2 + 4k - 4 = 4k - 2\) where \(1 \le k \le 10\)

OpenStudy (anonymous):

k thank you

OpenStudy (aum):

Assume \(k_1, k_2, k_3, k_4, k_5\) represents those k values whose terms add up to 100. Then, \(4(k_1+k_2+k_3+k_4+k_5) - 10 = 100 \\ 4(k_1+k_2+k_3+k_4+k_5) = 110 \\ (k_1+k_2+k_3+k_4+k_5) = 27.5 \) But \(k_1, k_2, k_3, k_4, k_5\) are all integers from 1 to 10. They cannot add up to 27.5. So there is no solution.

OpenStudy (aum):

In case the third line above is not clear, since the \(k^{th}\) term is \(4k - 2\), the \(k_1^{th}\) term is: \(4k_1 - 2\) \(k_2^{th}\) term is: \(4k_2 - 2\) etc...

OpenStudy (aum):

I don't know if the same number can be picked more than once.

OpenStudy (triciaal):

i don't have a solution but what I had was the average is 100/5 = 20 the numbers chosen deviate as -18, -14,-10,-6,-2, 2, 6, 10,14, 18 then as @aum states there is an arithmetic progression with d = 4

OpenStudy (aum):

I think tricia's method changes the problem to: pick 5 numbers from -18, -14,-10,-6,-2, 2, 6, 10,14, 18 that add up to zero.

OpenStudy (triciaal):

correction d =-4

OpenStudy (aum):

I think it is d = +4.

OpenStudy (triciaal):

I need to go to bed thanks add 4

OpenStudy (anonymous):

This is the first time I have attempted a "proof" on this site. Refer to the Mathematica calculation attached.